Question text
Calculate the minimum sample size for a study of the cost of 1 full year (tuition as well as room and board) when the level of confidence is 99%, the standard deviation is $1000 and the allowable error is $1600.

Enter your response as a whole number.

Answer by jim_thompson5910(35256)

ME = z*s/sqrt(n)


ME*sqrt(n) = z*s


sqrt(n) = z*s/ME


n = (z*s/ME)^2


n = (2.576*1000/1600)^2 ... z = 2.576 is found using a table (critical value for 99% confidence interval)


n = 2.5921


n = 3 ... round UP to the nearest whole number


The min sample size is 3